OpenStudy (anonymous):

What is the volume of the sphere with radius 4?

OpenStudy (imstuck):

Do you know the volume formula for a sphere?

OpenStudy (anonymous):

options a.64pi units^3 b.85.33pi units^3 c.256 pi units^3 d.5.33 pi units^3

OpenStudy (anonymous):

the formula is 4/3 pi r^3

OpenStudy (kainui):

\[\Large V=\frac{4}{3}\pi r^3\] This is the volume of a sphere from the radius. As an extra bonus, if you multiply this by the exponent and subtract one from the exponent you get the surface area of a sphere. \[\Large A= 4 \pi r^2\] It might be helpful to remember them as a pair.

OpenStudy (kainui):

So if you have the formula, just plug in the radius by replacing r with the value.

OpenStudy (nurali):

V=4/3πr^3 r=3 Plug the Value

OpenStudy (imstuck):

that's right. You have everything now except the volume itself. Use the radius they gave you of 4 and cube it. V = 4/3 pi (4)^3; V = 85.33 pi units cubed

OpenStudy (anonymous):

when i do the equation i get 268.0825

OpenStudy (anonymous):

@IMStuck

hartnn (hartnn):

keep pi as pi only. calculate 4/3 r^3 and append pi in the end